Grade 9Mechanics

If three particles of masses 2kg,1kg,3kg are placed at corners of an equilateral triangle of perimeter 6m then the distance of centre of mass of particles from 1kg mass is(assume 2kg on x axis)

Dheeraj
9 Years agoGrade 9

1 Answer

Shivansh Saumya
9 Years ago
let 2kg,3kg be on the x-axis
each side is 6m long.
co-ordinates of 1kg = (6,0)
co-ordinates of 2kg = (0,0)
co-ordinates of 3kg = (3,3*(31/2)) (pythagoras th.)
equation for co-ordinates of c.o.m.=
∑mixi/∑mi
∑miyi/∑mi
X = (1(6)+2(0)+3(3))/1+2+3 = 2.5
Y = (1(0)+2(0)+3(3*(3^1/2)))/1+2+3 = (3*(31/2))/2
co-ordinates = (2.5,(3*(31/2))/2)
Using distance formula = ((x2-x1)2+(y2-y1)2)1/2 
= 9 + 6.75 = 15.75 units